# TUS-REC2024: A Challenge to Reconstruct 3D Freehand Ultrasound Without External Tracker

### v1

@@ 4.6 Awards
Results from all participants were publicly displayed on the official leaderboards, except in cases where submissions encountered errors during the evaluation process. Additional certificates of recognition were awarded to the first-place team and the runner-up. All teams with successfully evaluated submissions received certificates of participation.
@@ 4.7 Timeline
The TUS-REC2024 Challenge is an open-call event designed to encourage broad community participation and, although this edition was structured as a one-time event tied to MICCAI 2024, its infrastructure and open-submission framework support potential future iterations, enabling continued engagement beyond the initial evaluation cycle.
The official timeline is aligned with MICCAI 2024, as detailed in Fig. 3. The Challenge began with the website launch and team registration on April 1, 2024, followed by the release of training data on May 13, and baseline code on June 23. On July 29, validation data and the Docker image template were released, offering participants a clear structure for submission. This template, along with an evaluation script that incorporates the Challenge metrics, aimed to ensure transparency and reproducibility in assessment and was designed to align with the BIAS Reporting Guideline. The submission window officially opened on August 12 and closed on September 9. The announcement of the winning teams took place on September 16, and the TUS-REC2024 Challenge event was held on October 6, 2024, during MICCAI 2024.
@@ 5.1 Participation Statistics
Fig. 4 presents the participant statistics for TUS-REC2024 Challenge. By the submission deadline, a total of 101 individuals registered, representing 43 teams comprising members from both academia and industry. Participants came from 14 countries across 5 continents, reflecting the international interest and global reach of the Challenge. Despite strong initial engagement, six teams submitted their Docker images by the submission deadline, involving a total of 25 participants. In total, 21 valid docker images were received. The number of submissions varied across teams, with several teams submitting multiple Docker images for performance optimisation. This decline from registration to submission may reflect the technical complexity of the task or limited preparation time. Notably, the majority of registered and submitting teams were affiliated with academic institutions, particularly universities. Overall, the statistics highlight both the broad appeal of the Challenge, and the practical hurdles faced by participants in progressing from initial registration to successful submission, such as time constrains and difficulties in model development.
@@ 5.2 Methods
The baseline method utilises the EfficientNet-B1 architecture [41], taking as input a pair of adjacent ultrasound frames. The network predicts a 6-DoF transformation, representing the transformation from the image coordinate system (in mm) of one frame to that of the other. The training loss is formulated as the mean squared error (MSE) between point coordinates transformed by the ground truth and those transformed by the network’s prediction.
@@
where $D(,\cdot,)$ denotes the MSE computed over the $x$, $y$ and $z$ coordinates of corresponding points. $T^{gt}_{j\leftarrow i}$ and $T_{j\leftarrow i}$ represent the ground truth and predicted transformation matrices, obtained by converting the predicted 6-DoF into a homogeneous transformation matrix. $T_{j\leftarrow i}$ is defined as in Eq. (1). $\mathbf{p}_{corner}$ denotes the coordinates of the four corner pixels in the image coordinate system (in pixels).
During inference, the 6-DoF between adjacent frames are estimated by sequentially inputting frame pairs into the network. The local DDFs for all pixels within an image are computed by left-multiplying the predicted local transformation matrix with the calibrated image coordinates and subtracting the coordinates of the reference frame: $DDF_{local}^{(i)}=T_{local}^{(i)}\cdot T_{scale}\cdot\mathbf{p}-T_{scale}\cdot\mathbf{p}$, where $\mathbf{p}$ denotes the coordinates of all pixels within an image in the image coordinate system (in pixels). $T_{local}^{(i)}=T_{i-1\leftarrow i}$ is the transformation matrix from frame $i$ to frame $i-1$, converted from the 6-DoF. The global transformation from any frame $i$ to the first frame, $T_{global}^{(i)}$, is derived by composing predicted local transformations through left-multiplication in reverse temporal order: $T_{global}^{(i)}=T_{local}^{(2)}\cdot T_{local}^{(3)}\cdots T_{local}^{(i)}$. The global DDFs for all pixels within an image are computed using $DDF_{global}^{(i)}=T_{global}^{(i)}\cdot T_{scale}\cdot\mathbf{p}-T_{scale}\cdot\mathbf{p}$. The local and global DDFs at predefined landmark locations can be obtained either by indexing the corresponding positions from $DDF_{local}$ and $DDF_{global}$, respectively, or by calculating them using the formula above, replacing all pixel coordinates $\mathbf{p}$ with the landmark locations.
5.2.1.2 FiMoNet
Fine-grained spatio-temporal learning is essential for freehand 3D ultrasound reconstruction. To address the complexities of long-range dependencies introduced by diverse probe motions as well as the large number of patches involved in spatio-temporal modeling, we adapted Mamba [42]. Mamba utilises the state-space model’s capacity to manage long-range dependencies, providing an effective solution for this task.
We employed ensemble learning to combine two models:
@@ 5.2.1.2 FiMoNet
For the 6-DoF transformations between adjacent frames, estimated by the network as $\theta$ and the ground truth 6-DoF transformation $\theta^{gt}$, we employed both the $L_{1}$ loss and Pearson correlation loss:
@@
The proposed approach utilises an EfficientNet-based CNN (pre-trained on ImageNet) to extract features from pairs of consecutive frames. These features are processed sequentially by a LSTM network to model temporal dependencies. The network predicts 3D translation and rotation parameters for each frame pair. Training is performed by minimising the MSE loss on these parameters, enabling the model to learn robust frame-to-frame alignments. TrivialAugment is used for data augmentation, and sequences of 16 frames are processed at a time, with adjacent frames concatenated prior to input into the CNN.
We employed a 5-fold cross-validation strategy for training; however, one fold displayed instability and was consequently excluded. The remaining four folds were ensembled by computing the median of the predicted 6-DoF transformation parameters, yielding a single final prediction. To further enhance the performance, a second 4-fold ensemble was trained on data downsampled by a factor of 1.25. The final submission is composed of both 4-fold ensembles (original and downsampled), resulting in a total of eight models.
During inference, the model estimates the rigid transformation parameters between each consecutive pair of frames within a scan. By sequentially concatenating these pairwise transformations starting from the first frame, we compute the global pose of each frame relative to the first frame. Using all frames’ global transformations, the 3D volume is reconstructed. The DDF is derived by back-mapping voxel coordinates from a reference 3D grid to their original frame positions. The model only explicitly predicts local transformations, while global transformations are obtained by sequentially accumulating these local estimates. Both local and global transformations are rigid and derived from the model’s frame-to-frame predictions.
@@ 5.2.1.4 FlowNet
The network is based on EfficientNet-B6, taking $n=10$ consecutive ultrasound frames $S\in\mathbb{R}^{n\times h\times w}$ as input. It outputs a set of transformation parameters $Y\in\mathbb{R}^{(n-1)\times 6}$, where each 6-DoF vector represents the rigid transformation from the last ultrasound frame $S^{(n)}$ to a preceding frame $S^{(i)},\,i\in[1,n-1]$. $Y$ is used to compute a flow field $F$, enabling the warping of $n-1$ frames to generate $S^{warp}\in\mathbb{R}^{(n-1)\times h\times w}$. $Y$ can also be expressed as matrices $T_{Y}$, where each $T_{Y}^{(i)}$ denotes the transformation matrix from $S^{(n)}$ to $S^{(i)}$. The transformation matrix between any two frames $S^{(i)}$ and $S^{(j)}$ can be obtained by calculating $T_{Y}^{(j)}\cdot(T_{Y}^{(i)})^{-1}$, forming the dense transformation matrix set $T^{ds}$. The resulting dense point coordinates $P^{ds}$ are then used to calculate the overall loss:
@@
Three models were selected: the final epoch model, the model from 100 epochs earlier, and the one with the lowest validation distance. Given the full scan $\mathcal{S}\in\mathbb{R}^{N\times h\times w}$, with local and global transformations $T_{local},T_{global}\in\mathbb{R}^{(N-1)\times 4\times 4}$, we sequentially process sequences of $n$ frames using a stride of $n-1$, such that the last frame of one sequence is the first of the next. Predictions from the three models are averaged to obtain the final $Y$ and $T_{Y}$. For the first sequence, local transformation could be calculated by $T_{local}^{(i)}=T_{Y}^{(i-1)}\cdot(T_{Y}^{(i)})^{-1}$, and global transformation is calculated by $T_{global}^{(i)}=T_{Y}^{(1)}\cdot(T_{Y}^{(i)})^{-1}$. After computing local transformations for all frames in the first sequence, subsequent sequences are processed sequentially using the same method to obtain local transformations. The global transformation for the $k^{th}$ frame in $l^{th}$ sequence $S^{((l-1)\times(n-1)+k)}$ is computed as $T_{global}^{((l-1)\times(n-1)+k)}=T^{(1)}_{Y,s_{1}}\cdot T^{(1)}_{Y,s_{2}}\cdots T^{(1)}_{Y,s_{l}}\cdot(T_{Y,s_{l}}^{(k)})^{-1}$, where $T^{(k)}_{Y,s_{l}}$ denotes the transformation matrix from $n^{th}$ frame to the $k^{th}$ frame in the $l^{th}$ sequence. This yields $T_{local}$ and $T_{global}$ for the full scan. The scan is then reversed, and the same procedure is applied to obtain $T_{local}^{reverse}$ and $T_{global}^{reverse}$. Final predictions are obtained by averaging the forward and reversed results: $T^{avg}_{local}=(T_{local}+T_{local}^{reverse})/2$, $T^{avg}_{global}=(T_{global}+T_{global}^{reverse})/2$.
@@
5.2.1.5 MoGLo-Net [^14] https://github.com/guhong3648/US3D
Input images are cropped to square regions to remove background artifacts and normalised to $[-1,1]$. As shown in Fig. 5, we developed MoGLo-Net, a motion-based learning network with global-local attention. Two ultrasound sequences, each consisting of $k+1$ frames, are processed in parallel through a ResNet-based encoder for consistent feature refinement.
@@
The encoded features are merged into $M$ and refined with $C$ via encoder blocks (purple arrows). The resulting features are concatenated to form the final feature map $F$, where the global-local attention module is applied. This module is designed as a self-attention mechanism [46, 47, 48]: (1) global (GF) and local features (LF) are downsampled by factors of 4 and 2, respectively; (2) LFs are extracted from early encoder layers as patch-wise feature blocks; (3) both GFs and LFs are refined using a conventional attention mechanism; (4) cosine similarity between GFs and LFs serves as attention scores to weight LFs; and (5) weighted LFs are projected to aggregate local information. The final recalibrated feature, formed by concatenating GFs and LFs, is fed into the RNN-based estimator (Conv-GRU) to predict 6-DoF.
We employed three loss functions: Correlation Loss [11] ensures motion consistency; Triplet Loss [11] contrasts the final feature maps; and Motion-based Mean Absolute Error (MMAE) emphasises errors in fast-motion regions:
@@
where $s+1$ denotes the number of frames in an ultrasound sequence. $w_{i}=\left|\Delta{\theta}_{i}\right|+\varepsilon$ is a motion-based weighting term. Fast motion errors are penalised more heavily, as $w_{i}$ increases with larger motion vectors. The smoothing term $\varepsilon$ prevents over-amplification. Our model predicts 6-DoF for ultrasound sequences, but only the final frame’s estimation is used during inference. Therefore, predicting the scan motion for the entire scan requires $N$ sequences, where $N$ is the number of frames in the scan. To do this, we add padding at the beginning of the scan frames. The global and local transformations are derived based on the TUS-REC2024 baseline code.
@@
To address the complexities in freehand ultrasound reconstruction, particularly out-of-plane motion [49], we proposed a lightweight, physics-informed deep learning model. Our dual-stream network decouples spatial and temporal learning, incorporating learnable operators to capture data priors for modeling temporal relationships and integrating a physical model to simplify learning, offering flexibility for various scanning paths.
The PLPPI model consists of spatial and temporal branches, followed by a fusion module and prediction head. The spatial branch uses 2D convolutions to aggregate intra-frame spatial context, while the temporal branch extracts inter-frame motion cues via speckle decorrelation patterns. This involves constructing a correlation volume to quantify the underlying motion information. The outputs from both branches are fused to represent the input sequence in the feature space, with speckle decorrelation serving as a key physics-based prior. The temporal branch computes a correlation volume $cv$ by measuring patch-wise similarity between two dense feature maps, $c_{21}$ and $c_{22}$, where $c\in\mathbb{R}^{h\times w\times d}$. $cv$ is defined as: $cv(x_{1},x_{2})=\sum_{s\in[-p,p]\times[-p,p]}c_{21}(x_{1}+s)^{T}\cdot c_{22}(x_{2}+s)$, with $x_{1}$, $x_{2}$ denoting the patch locations centered at $c_{21}$ and $c_{22}$, respectively, and $p$ the maximum displacement between $x_{1}$ and $x_{2}$. The squared patch size is $K=2p+1=21$. During training, input image stacks are split into two sub-volumes and passed through 2D convolutions to obtain $c_{21}$ and $c_{22}$. Temporal features are then bilinearly upsampled and fused with spatial features for joint representation.
Compared to our previous approach [49], we introduced two key modifications: (1) replacing the ResNet backbone with the pretrained foundation model Biomedical CLIP [50], and (2) redesigning the loss function to better leverage the capabilities of the foundation model. The new loss is defined as:
@@
During inference, the model outputs $n-1$ local transformations estimated from $n$ input images. Sliding window averaging is applied to obtain the final local transfromations: ${\theta}_{local}^{(i)}=\frac{1}{W}\sum^{i}_{j=i-W+1}\theta^{(j)}$, with window size $W$. Global transformation is then computed as $T_{global}^{(i)}=\prod^{i}_{j=1}T_{local}^{(j)}$ where $T_{local}^{(j)}$ is converted from ${\theta}_{local}^{(j)}$.

### v2

@@ 4.6 Awards
Results from all participants were publicly displayed on the official leaderboard [^9] https://github-pages.ucl.ac.uk/tus-rec-challenge/TUS-REC2024/leaderboard.html, except in cases where submissions encountered errors during the evaluation process. Additional certificates of recognition were awarded to the first-place team and the runner-up. All teams with successfully evaluated submissions received certificates of participation. Members of the organizers’ institutes may participate but not eligible for awards and not listed in leaderboard.
@@ 4.7 Timeline
The TUS-REC2024 Challenge is an open-call event designed to encourage broad community participation. Although this edition was structured as a one-time event tied to MICCAI 2024, its infrastructure and open-submission framework support potential future iterations, enabling continued engagement beyond the initial evaluation cycle.
The official timeline is aligned with MICCAI 2024, as detailed in Fig. 4. The Challenge began with the launch of website and team registration on April 1, 2024, followed by the release of training data on May 13, and release of baseline code on June 23. On July 29, validation data and the Docker image template were released, offering participants a clear instruction for submission. This template, along with an evaluation script that incorporates the Challenge metrics, aimed to ensure transparency and reproducibility in assessment and was designed to align with the BIAS Reporting Guideline. The submission window officially opened on August 12 and closed on September 9. The announcement of the winning teams took place on September 16, and the TUS-REC2024 Challenge event was held on October 6, 2024, during MICCAI 2024. The participating teams can publish their results separately after publication of the joint challenge paper.
@@ 5.1 Participation Statistics
Fig. 5 presents the participant statistics for TUS-REC2024 Challenge. By the submission deadline, a total of 101 individuals registered, representing 43 teams comprising members from both academia and industry. Participants came from 14 countries across 5 continents, reflecting the international interest and global reach of the Challenge. Despite strong initial engagement, 6 teams submitted their Docker images by the submission deadline, involving a total of 25 participants. In total, 21 valid docker images were received. The number of submissions varied across teams, with several teams submitting multiple Docker images for performance optimisation. This decline from registration to submission may reflect the technical complexity of the task or limited preparation time. Notably, the majority of registered and submitting teams were affiliated with academic institutions, particularly universities. Overall, the statistics highlight both the broad appeal of the Challenge, and the practical hurdles faced by participants in progressing from initial registration to successful submission, such as time constraints and difficulties in model development.
@@ 5.2 Methods
The baseline method utilises the EfficientNet-B1 architecture [86], taking as input a pair of adjacent ultrasound frames. The network predicts a 6-DoF transformation, representing the transformation from the image coordinate system (in mm) of one frame to that of the other. The training loss is formulated as the mean squared error (MSE) between point coordinates transformed by the ground truth and the predicted transformations:
@@
where $D(,\cdot,)$ denotes the MSE computed over the $x$, $y$ and $z$ coordinates of corresponding points. $T_{j\leftarrow i}$ is obtained by converting the predicted 6-DoF into a homogeneous transformation matrix. $T^{gt}_{j\leftarrow i}$ represents the ground truth transformation matrix, as defined in Eq. (1). $\mathbf{p}_{corner}$ denotes the pixel coordinates of the four corner points in the image coordinate system.
During inference, the 6-DoF transformations between adjacent frames are estimated by sequentially inputting frame pairs into the network. The local DDFs are computed by $DDF_{local}^{(i)}=T_{local}^{(i)}\cdot T_{scale}\cdot\mathbf{p}-T_{scale}\cdot\mathbf{p}$, where $\mathbf{p}$ denotes the coordinates of all pixels within an image in the image coordinate system (in pixels). $T_{local}^{(i)}=T_{i-1\leftarrow i}$ is the transformation matrix from frame $i$ to frame i1i-1, converted from the 6-DoF. The global transformation from any frame $i$ to the first frame, $T_{global}^{(i)}$, is derived by composing predicted local transformations through $T_{global}^{(i)}=T_{local}^{(2)}\cdot T_{local}^{(3)}\cdots T_{local}^{(i)}$. The global DDFs are computed using $DDF_{global}^{(i)}=T_{global}^{(i)}\cdot T_{scale}\cdot\mathbf{p}-T_{scale}\cdot\mathbf{p}$. The local and global DDFs at predefined landmark locations can be obtained either by indexing the corresponding positions from $DDF_{local}$ and $DDF_{global}$, respectively, or by calculating them using the formula above, replacing all pixel coordinates $\mathbf{p}$ with the landmark locations.
5.2.1.2 FiMoNet [^14] https://github.com/Lmy0217/FiMoNet
Fine-grained spatio-temporal learning is essential for trackerless freehand 3D ultrasound reconstruction. FiMoNet adapts Mamba [52] to address the complexities of long-range dependencies introduced by diverse probe motions as well as the large number of patches involved in spatio-temporal modeling. Mamba utilises the state space model’s capacity to manage long-range dependencies, providing an effective solution for this task.
Ensemble learning is used to combine two models:
@@
For the 6-DoF transformations between adjacent frames, estimated by the network as $\theta$ and the ground truth 6-DoF transformation $\theta^{gt}$, both the $L1$ loss and Pearson correlation loss are used:
@@
This approach utilises an EfficientNet-based CNN (pre-trained on ImageNet) to extract features from pairs of consecutive frames. These features are processed sequentially by a LSTM network to model temporal dependencies. The network predicts 3D translation and rotation parameters for each frame pair. Training is performed by minimising the MSE loss on these parameters, enabling the model to learn robust frame-to-frame alignments. TrivialAugment [87] is used for data augmentation, and sequences of 16 frames are processed at a time, with adjacent frames concatenated prior to input into the CNN.
RecuVol applied a 5-fold cross-validation strategy for training. However, one fold displayed instability and was consequently excluded. The remaining four folds were ensembled by computing the median of the predicted 6-DoF transformation parameters, yielding a single final prediction. To further enhance the performance, a second 4-fold ensemble was trained on data downsampled by a factor of 1.25. The final submission was composed of both 4-fold ensembles (original and downsampled), resulting in a total of eight models.
During inference, the model estimates the rigid transformation parameters between each consecutive pair of frames within a scan. By sequentially concatenating these pairwise transformations starting from the first frame, the method computes the global pose of each frame relative to the first frame. Using all frames’ global transformations, the 3D volume is reconstructed. The DDF is derived by back-mapping voxel coordinates from a reference 3D grid to their original frame positions. The model only explicitly predicts local transformations, while global transformations are obtained by sequentially accumulating these local estimates. Both local and global transformations are rigid and derived from the model’s frame-to-frame predictions.
@@ 5.2.1.4 FlowNet
The network is based on EfficientNet-B6, taking $n=10$ consecutive ultrasound frames $S\in\mathbb{R}^{n\times h\times w}$ as input. It outputs a set of transformation parameters $Y\in\mathbb{R}^{(n-1)\times 6}$, where each 6-DoF vector represents the rigid transformation from the last ultrasound frame $S^{(n)}$ to a preceding frame $S^{(i)},\,i\in[1,n-1]$. $Y$ is used to compute a flow field $F$, enabling the warping of n1n-1 frames to generate $S^{warp}\in\mathbb{R}^{(n-1)\times h\times w}$. $Y$ can also be converted to matrices $T_{Y}$, where each $T_{Y}^{(i)}$ denotes the transformation matrix from $S^{(n)}$ to $S^{(i)}$. The transformation matrix between any two frames $S^{(i)}$ and $S^{(j)}$ can be obtained by calculating $T_{Y}^{(j)}\cdot(T_{Y}^{(i)})^{-1}$, forming the dense transformation matrix set $T^{ds}$. The resulting dense point coordinates $P^{ds}$ are then used to calculate the overall loss:
@@
Given the full scan $\mathcal{S}\in\mathbb{R}^{N\times h\times w}$, with local and global transformations $T_{local},T_{global}\in\mathbb{R}^{(N-1)\times 4\times 4}$, sequences of $n$ frames are sequentially processed using a stride of n1n-1, such that the last frame of one sequence is the first of the next. Three models are selected: the final epoch model, the model from 100 epochs earlier, and the one with the lowest validation distance. Predictions from the three models are averaged to obtain the final $Y$ and $T_{Y}$. For the first sequence, local transformation could be calculated by $T_{local}^{(i)}=T_{Y}^{(i-1)}\cdot(T_{Y}^{(i)})^{-1}$, and global transformation is calculated by $T_{global}^{(i)}=T_{Y}^{(1)}\cdot(T_{Y}^{(i)})^{-1}$. After computing local transformations for all frames in the first sequence, subsequent sequences are processed sequentially using the same method to obtain local transformations. The global transformation for the $k^{th}$ frame in $l^{th}$ sequence $S^{((l-1)\times(n-1)+k)}$ is computed as $T_{global}^{((l-1)\times(n-1)+k)}=T^{(1)}_{Y,s_{1}}\cdot T^{(1)}_{Y,s_{2}}\cdots T^{(1)}_{Y,s_{l}}\cdot(T_{Y,s_{l}}^{(k)})^{-1}$, where $T^{(k)}_{Y,s_{l}}$ denotes the transformation matrix from $n^{th}$ frame to the $k^{th}$ frame in the $l^{th}$ sequence. This yields $T_{local}$ and $T_{global}$ for the full scan. The scan is then reversed, and the same procedure is applied to obtain $T_{local}^{reverse}$ and $T_{global}^{reverse}$. Final predictions are obtained by averaging the forward and reversed results: $T^{avg}_{local}=(T_{local}+T_{local}^{reverse})/2$, $T^{avg}_{global}=(T_{global}+T_{global}^{reverse})/2$.
@@
5.2.1.5 MoGLo-Net [^18] https://github.com/guhong3648/MoGLo
Input images are cropped to square regions to remove background artifacts and normalised to [1,1][-1,1]. Fig. 6 shows overview of MoGLo-Net, a motion-based learning network with global-local attention. Two ultrasound sequences, each consisting of $s+1$ frames, are processed in parallel through a ResNet-based encoder for consistent feature refinement.
@@
The encoded features of the two input sequences are merged into $M$ and refined via encoder blocks (purple arrows) together with $C$. The resulting features are concatenated to form the final feature map $F$, where the global-local attention module is applied. This module is designed as a self-attention mechanism [89, 90, 91]: (1) global (GF) and local features (LF) are downsampled by factors of 4 and 2, respectively; (2) LFs are extracted from early encoder layers as patch-wise feature blocks; (3) both GFs and LFs are refined using a conventional attention mechanism; (4) cosine similarity between GFs and LFs serves as attention scores to weight LFs; and (5) weighted LFs are projected to aggregate local information. The final recalibrated feature, formed by concatenating GFs and LFs, is fed into the RNN-based estimator (Conv-GRU) to predict 6-DoF transformations $\Delta\theta$.
Three loss functions are used: Correlation Loss [38] ensures motion consistency; Triplet Loss [38] contrasts the final feature maps; and Motion-based Mean Absolute Error (MMAE) emphasises errors in fast-motion regions:
@@
where $s+1$ is the length of input sequence and $k$ denotes dimensions of 6-DoF prediction. $w_{i}=\left|\Delta{\theta}_{i}\right|+\varepsilon$ is a motion-based weighting term. Fast motion errors are penalised more heavily, as $w_{i}$ increases with larger motion vectors. The smoothing term $\varepsilon$ prevents over-amplification. This model predicts 6-DoF for the input sequences, but only the final frame’s estimation is used during inference. Therefore, predicting the motion for the entire scan requires $N$ sequences, where $N$ is the number of frames in the whole scan. To do this, we add padding at the beginning of the scan frames. The global and local transformations are derived based on the TUS-REC2024 baseline code.
@@
To address the complexities in trackerless freehand ultrasound reconstruction, particularly out-of-plane motion [62], a lightweight, physics-informed deep learning model is proposed. The dual-stream network decouples spatial and temporal learning, incorporating learnable operators to capture data priors for modeling temporal relationships and integrating a physical model to simplify learning, offering flexibility for various scanning paths.
The PLPPI model consists of spatial and temporal branches, followed by a fusion module and prediction head. The spatial branch uses 2D convolutions to aggregate intra-frame spatial context, while the temporal branch extracts inter-frame motion cues via speckle decorrelation patterns. This involves constructing a correlation volume to quantify the underlying motion information. The outputs from both branches are fused to represent the input sequence in the feature space, with speckle decorrelation serving as a key physics-based prior. During training, input image stacks are split into two sub-volumes and passed through 2D convolutions to obtain $c_{21}$ and $c_{22}$. The temporal branch computes a correlation volume $cv$ by measuring patch-wise similarity between $c_{21}$ and $c_{22}$, where $c\in\mathbb{R}^{h\times w\times d}$. $cv$ is defined as: $cv(x_{1},x_{2})=\sum_{s\in[-p,p]\times[-p,p]}c_{21}(x_{1}+s)^{T}\cdot c_{22}(x_{2}+s)$, with $x_{1}$, $x_{2}$ denoting the patch locations centered at $c_{21}$ and $c_{22}$, respectively, and $p$ the maximum displacement between $x_{1}$ and $x_{2}$. The squared patch size is $K=2p+1=21$. Temporal features are then bilinearly upsampled and fused with spatial features for joint representation.
Compared to their preceding work [62], two key modifications are introduced: (1) replacing the ResNet backbone with the pretrained foundation model Biomedical CLIP [92], and (2) redesigning the loss function to better leverage the capabilities of the foundation model. The new loss is defined as:
@@
During inference, the model outputs n1n-1 local transformations estimated from $n$ input images. Sliding window averaging is applied to obtain the final local transfromations: ${\theta}_{local}^{(i)}=\frac{1}{W}\sum^{i}_{j=i-W+1}\theta^{(j)}$, with window size $W$. Global transformation is then computed as $T_{global}^{(i)}=\prod^{i}_{j=1}T_{local}^{(j)}$ where $T_{local}^{(j)}$ is converted from ${\theta}_{local}^{(j)}$.